Problem Solving


1. A television station broadcasts in the frequency range of 470 to 478 MHz.

Calculate the corresponding wavelength range. Is it a VHF or a UHF station?


2. Find the frequencies of the following:

a. FM radio wave with a wavelength of 3 m

b. microwave with a wavelength of 0.03 m

c. x-rays with a wavelength 0f 0.5 mm.


3. If the frequency of an EM wave is 106 Hz, what is its wavelength in

nanometers? in meters? What type of electromagnetic wave is this?


1
Expert's answer
2021-01-23T10:30:20-0500

(1) "\lambda=c/\nu"


"\lambda_1=c/\nu_1=3\cdot10^8/478\cdot10^6=0.627(m)"


"\lambda_2=c/\nu_2=3\cdot10^8/470\cdot10^6=0.638(m)"


UHF station.


(2) "\nu=c/\lambda"


a) "\nu=3\cdot10^8/3=1\cdot10^8(Hz)"


b) "\nu=3\cdot10^8/0.03=1\cdot10^{10}(Hz)"


c) "\nu=3\cdot10^8/0.5\cdot10^{-3}=6\cdot10^{11}(Hz)"


(3) "\lambda=c/\nu=3\cdot10^8/106=2.83\cdot10^6(m)=2.83\cdot10^{15}(nm)"


Radio wave (SLF)